OpenStudy (anonymous):

Eaton purchased a boat for $11,975. He made a down payment of $3,500. He applied for a five-year installment loan with an interest rate of 11.2%. What is the total cost of the boat after five years? $13,316.20 $14,606.60 $15,693.69 $11,106.81

OpenStudy (ivettef365):

principal amount would be 11975 - 3500

OpenStudy (amistre64):

\[B_n=B_ok^{c}-P\frac{1-k^c}{1-k}\] assuming we want a zero baolance after 5 years; and i assume monthly payments \[0=B_ok^{12*5}-P\frac{1-k^{12*5}}{1-k}~:~k=1+\frac r{12}\] \[B_ok^{12*5}=P\frac{1-k^{12*5}}{1-k}\] \[B_ok^{12*5}\frac{1-k}{1-k^{12*5}}=P\] 5 years worth of payments, minus the intial amount; gives us total cost loan

OpenStudy (amistre64):

well, i spose if we just add up the payments itll be fine ... i was thinking total interest cost on the loan

OpenStudy (anonymous):

I'm not familiar with that equation, what does everything stand for?

OpenStudy (ivettef365):

did you try the website

OpenStudy (amistre64):

Bo is initial balance (loan amount) k is a clean up variable to represent that compounding stuff and P is the Payments

OpenStudy (anonymous):

I don't really like using website calculators, they're often wrong or make an error @ivettef365

OpenStudy (amistre64):

assuming monthly payments of 185.12 ... and there are 60 of them in 5 years

OpenStudy (anonymous):

Can you plug everything in for me? I'm still confused

OpenStudy (amistre64):

what is your inital loan amount?

OpenStudy (anonymous):

8475?

OpenStudy (amistre64):

good, and what is your compounding setup? 1 + r/12

OpenStudy (anonymous):

1 + 0.112/ 12

OpenStudy (amistre64):

good, and putting it all together we get P = 8475k^(12*5)(1-k)/(1-k^(12*5)) since there are 12*5 payments .. 60 payments 60(8475k^(12*5))(1-k)/(1-k^(12*5)), k=1+.112/12 http://www.wolframalpha.com/input/?i=60%288475k%5E%2812*5%29%29%281-k%29%2F%281-k%5E%2812*5%29%29%2C+k%3D1%2B.112%2F12

OpenStudy (anonymous):

Aaaahhh I'm still confused

OpenStudy (anonymous):

Idk how to plug this into my calculator

OpenStudy (amistre64):

60 * 8475 * (1+.112/12)^(60) * (1-(1+.112/12)) / (1-(1+.112/12)^(60))

OpenStudy (amistre64):

might help to approximate that k stuff

OpenStudy (dumbcow):

yeah formula for solving for monthly payments is a bit complex...what class is this for? you would be best to use a financial calculator total payments = 60*monthly payment

OpenStudy (ivettef365):

you could off have the answer using website, I got the right answer

OpenStudy (dumbcow):

yeah but cant use website on test or in class....get a financial calculator like a TI BAII

OpenStudy (anonymous):

I have a TI-36X Pro

OpenStudy (anonymous):

@amistre64 I plugged everything into my calculator and got 11,106.81

OpenStudy (dumbcow):

@murrcat , that is correct

OpenStudy (anonymous):

Sweet

OpenStudy (amistre64):

i developed my own formula .... which is a less "simplified" version of most texts i think

OpenStudy (anonymous):

11,106.81 was wrong :c

OpenStudy (amistre64):

you know what, we forgot to add in the downpayment as part of the cost ....

OpenStudy (amistre64):

11,100 3,500 ------- 14,600